OpenStudy (anonymous):

write an equation in slope form for the given line through the given point with the given slope. (-5,6); m=0.5

OpenStudy (anonymous):

Use the form y = mx + b y = 0.5x + b Just plug in the point. 6 = 0.5(-5) + b 6 = -2.5 + b b = 8.5 Then plug it back in. y = 0.5x + 8.5

OpenStudy (anonymous):

A. y+5=0.5 (x-6) B. y+6=0.5 (x-5) C. y+6=0.5 (x+5) D. y-6=0.5 (x+5)

OpenStudy (anonymous):

It would be D. I thought you meant slope-intercept form. This is point-slope form which is: \[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You're welcome